# The Host Galaxies of Fast Radio Bursts Track a Combination of Stellar Mass and Star Formation, Similar to Type Ia Supernovae

### v1

@@ Abstract
We develop a new statistical framework for studying the host galaxies of astrophysical sources that accounts for both redshift evolution and the multi-variate nature of host-galaxy properties. These aspects are critical when dealing with sources that span a wide range of redshifts, and/or with unknown redshift-dependent selection effects. We apply our method to a sample of Fast Radio Burst (FRB) host-galaxies as a means of probing the uncertain progenitor(s) of these events. Using our method we are able to rule out that FRBs track star-formation rate (SFR), as would be expected if FRBs are associated exclusively with young neutron stars born via core-collapse supernovae (SNe). Furthermore, we robustly rule out a recently proposed model whereby FRBs track SFR only above a certain metallicity threshold. Motivated by the fact that at least one FRB has been localized to a globular cluster (GC), we also investigate the hypothesis that that all FRBs occur in GCs and rule out this scenario explicitly for the first time. Alternatively, we find that a ‘mixed’ model whereby FRBs track a linear combination of both SFR and stellar-mass best explains the data. The preferred parameters of such a mixed model are remarkably similar to those inferred for Type Ia SNe, and implies a possible connection between the progenitors of these different transients.
@@
software: numpy (C. R. Harris et al., 2020) matplotlib (J. D. Hunter, 2007), emcee (D. Foreman-Mackey et al., 2013), astropy (Astropy Collaboration et al., 2013), scipy (P. Virtanen et al., 2020), pandas (Wes McKinney, 2010), statsmodels (S. Seabold & J. Perktold, 2010),
@@ 1 Introduction
One challenge to the naive magnetar scenario is the environment of some FRB sources. With the growing sample of localized FRBs whose host galaxies have been identified, it has been found that some FRBs reside in quiescent galaxies with very little recent star-formation (K. W. Bannister et al., 2019; V. Ravi et al., 2019; L. Connor et al., 2023; C. J. Law et al., 2024; V. Shah et al., 2025; T. Eftekhari et al., 2025). This is in tension with the conventional expectation for magnetars—that these sources should be formed through the core-collapse of massive stars. In this scenario, FRB sources should closely-track star-formation and their occurrence rate in quiescent galaxies should be fairly low. More direct evidence along this line has recently surfaced with the localization of a nearby FRB to a globular cluster (GC; F. Kirsten et al. 2022). Redshift zero GCs are characterized by old stellar populations with no ongoing star-formation, directly challenging standard magnetar formation channels. In light of these observations alternative formation channels for magnetar FRB progenitors have been proposed, including the possibility that magnetars could be formed in the accretion-induced-collapse (AIC) of a white dwarf (WD), or in the aftermath of a subset of binary NS mergers (B. Margalit et al., 2019; K. Kremer et al., 2021; W. Lu et al., 2022; T. Eftekhari et al., 2025). Questions pertaining to the progenitors of FRBs, and whether some or all of them are related to magnetars (and if so, what are the formation channels and characteristics of such magnetars) are therefore central to our understanding of FRB phenomena.
@@
A particularly powerful tool in addressing these questions is studying the population of host-galaxies and local environments of FRB sources. Although host-galaxy studies in general offer only circumstantial evidence about the nature of sources, the ability to statistically study large samples via this approach makes this method appealing. Indeed, host-galaxy studies have been central to building an understanding of other transient phenomena. They have helped solidify the association of (at least typical) long gamma-ray bursts (GRBs) with low-metallicity star-forming environments, consistent with expectations from the collapsar model (A. S. Fruchter et al., 2006); they have shown that superluminous supernovae (SLSNe) occur in similar environments to long GRBs, potentially consistent with rapidly-rotating magnetar progenitors (R. Lunnan et al., 2014); and they have shown that the formation channels of short GRB and Type Ia supernovae (SNe) progenitors differ significantly from other transients, tracking instead a combination of both star-formation and stellar mass (e.g., E. Scannapieco & L. Bildsten, 2005; W. Fong et al., 2013).
@@ 2 Past Work
In contrast, host-galaxy properties are intrinsically multivariate and characterized by a combination of several parameters, such as $M$, SFR, and redshift (at the very least). To use KS or AD tests, host-galaxy samples are therefore typically “marginalized” over a number of variables so that the distribution of some single variable (say stellar mass) may be examined. This approach, though useful, is not statistically sound.
For example, it is possible for two intrinsically different galaxy distributions in the $(M,{\rm SFR})$ plane to have similar 1D distributions in $M$ and/or SFR. Univariate KS or AD tests might therefore suggest that the two samples are drawn from the same distribution, when in fact they are not. One way to conceptualize this ambiguity is to consider that the choice of axes along which to marginalize is arbitrary in a multivariate phase-space. In particular, there is no apriori reason why KS tests should be run on the masses and SFRs of galaxies rather than on some alternative linear combination of the two. The fact that this arbitrary choice of axes will affect the KS score and, perhaps, the resulting conclusion is therefore troubling. Unfortunately, there is no distribution-free generalization of the KS or AD tests to more than one dimension (although see e.g., J. A. Peacock 1983 and G. Fasano & A. Franceschini 1987 for KS generalizations that are only weakly-dependent on the assumed distribution).111‘Distribution-free’ means that there is no statistical method to test whether two multivariate samples are drawn from the same underlying distribution, without making any assumptions about the functional form of the underlying distribution.
@@
This was first pointed out in the context of FRBs by C. D. Bochenek et al. (2021). In that work, the authors devised a method of ‘transposing’ host galaxies to redshift $z=0$. They then compared the transposed host galaxy distribution with a complete sample of core-collapse SN (CCSN) hosts at $z=0$ and found that the two samples are consistent with one another, but only after accounting for this redshift scaling (transposition). This illustrated the importance of accounting for redshift when making such comparisons, which can affect the overall conclusions markedly. However the particular methodology presented by C. D. Bochenek et al. (2021)—though novel and important—has one noticeable shortcoming. In that work, the stellar mass and SFR of each FRB host-galaxy were rescaled to $z=0$ based on the star-forming main sequence of galaxies. Specifically, each galaxy’s stellar-mass and SFR were transposed in a way that preserves the galaxy’s percentiles in both $M$ and SFR relative to the distribution of star-forming galaxies at the redshift of the transient and at $z=0$. Implicit in this choice is the assumption that FRB hosts evolve according to the redshift evolution of star-forming galaxies. However, there is no guarantee apriori that this be the case. Indeed, if one is interested in testing the hypothesis that FRBs track anything other than SFR, this choice may bias the results.
@@ 3.1 Extending the Distribution Function
Being trained on galaxy catalogs, the J. Leja et al. (2022) distribution function is limited by these catalog’s range and completeness. In particular the distribution is valid between redshifts $30.2<z<$ 3, and is only complete above a redshift-dependent mass-completeness limit $M_{\rm c}(z)$. The latter is due to lower mass galaxies being fainter and harder to detect in surveys. Of course, even though galaxy catalogs may not be complete to such galaxies, galaxies of mass $M<M_{\rm c}(z)$ are in reality abundant. In the FRB host-galaxy sample that we adopt there are two sources (out of 51) that fall below the threshold mass described above. To take these two galaxies into account, and to allow our formalism to be applied more broadly without limitation on galaxy masses, we extend the J. Leja et al. (2022) distribution below $M_{\rm c}$ using the following two considerations.
@@
Here $b(z)$ is the (redshift-dependent) low-mass slope of the star-forming sequence as given by J. Leja et al. (2022),222 The parameter $b$ is defined in Equation 9 of that paper. Its dependence on redshift is parameterized via their Equation 10, with best-fit coefficients listed in their Table 1. Here we adopt parameters as fit to the ‘ridge’ of the star-forming sequence. and $\Phi(M,z)$ is a double Schechter-function fit to the stellar mass function taken from J. Leja et al. (2022). Finally, we take $M_{\rm th}=2M_{\rm c}(z)$ where $M_{\rm c}(z)$ is the mass-completeness cutoff described in J. Leja et al. (2022) (e.g., their Table 1). We take a threshold mass that is a factor of two larger than $M_{\rm c}$ to avoid possible edge-effects that may affect the distribution near the mass-completeness cutoff.
@@
The extension described in Equation (3.1) effectively assumes that low mass galaxies with $M<M_{\rm th}$ follow the star-forming sequence (or more accurately, that they can be transposed with relation to the star-forming sequence, even if they fall far from its locus). In general, this is a strong assumption. Many galaxies are quiescent or transitioning and do not follow the star-forming sequence. Given that we are interested in testing the hypothesis that FRB hosts track SFR and/or stellar mass, it is crucial to account for these quiescent galaxies. Indeed, our present work is motivated by shortcomings of previous studies in this respect. Ideally we would not need to extend the distribution function in this way. However we must define some extension if we wish to include the two FRB hosts that fall below the mass-completeness threshold. We argue that Equation (3.1) is nevertheless a reasonable choice for the following reasons. First, low-mass galaxies—those that are affected by our extension—are very strongly biased towards being star-forming, while quiescent galaxies (which do not track the star-forming sequence) mostly have $M>M_{\rm th}$. In other words, using the star-forming sequence is a reasonable approximation for low mass galaxies, exactly the regime where Equation (3.1) is relevant. Second, there are only two sources in our FRB sample that are affected by this extension, and both lie only slightly below the cutoff mass. That is, the extrapolation introduced by our extension is very small, and only applies to a small fraction of our sample (see Table 2).
@@
One final extension we apply to the J. Leja et al. (2022) distribution function is a treatment that incorporates low-redshift $z<0.2$ galaxies, where the J. Leja et al. (2022) distribution is truncated. Ideally, a model that is trained on low-redshift galaxy catalogs would be incorporated within this framework and allow a self-consistent evolution with redshift. However this is well beyond the scope of our present work. In order to include low-redshift FRB host galaxies in our current analysis we therefore extend the J. Leja et al. (2022) distribution in redshift-space using the following approximation: we assume that the galaxy distribution evolves very little between $z=0.2$ and $z=0$, and therefore that the distribution function is constant at these redshifts. In practice, this implies the following extension,
[…]
@@
In Equation (2) we take $z=0.25$ as a cutoff redshift instead of $z=0.2$ in order to avoid possible edge-effects of the distribution near the minimal redshift.
@@
Our redshift extension may be overly simplified, however it is born out of necessity given that we wish to consider FRB host galaxies that would otherwise be excluded from our analysis given the redshift cutoff. Overall, our approach to extending the J. Leja et al. (2022) distribution function to low redshifts and low galaxy masses is very similar to the approach recently taken by K. Sharma et al. (2024) and N. Loudas et al. (2025). This is reassuring given that we developed our analysis independently of these works. One minor difference is that both K. Sharma et al. (2024) and N. Loudas et al. (2025) use the extrapolated $z<0.2$ redshift dependence implied by fitting functions to the star-forming sequence and the stellar-mass function obtained by J. Leja et al. (2019, 2022). These are extrapolations because the fits were obtained on data at $z>0.2$. In our work we instead assume that the galaxy distribution at $z<0.25$ is similar to that at $z=0.25$. This effectively neglects evolution at very low redshifts.
We note that both approaches have drawbacks, and the only self-consistent approach would be to train the model on data that goes down to $z=0$. Supplementary Figure 10 in K. Sharma et al. (2024) compares these two different approaches, and shows that they result in nearly identical stellar-mass distributions. On the other hand, we expect that the SFR distributions would show larger discrepancies between the two approaches. This is because the star-formation drops at low redshifts. Specifically, the cosmic star-formation density $\Psi(z)$ is $\simeq 0.26\,{\rm dex}$ lower at $z=0$ compared to $z=0.25$ (P. Madau & M. Dickinson, 2014). This suggests that switching to the methodology of K. Sharma et al. (2024) and N. Loudas et al. (2025) would systematically shift our theoretical SFR distributions downwards by $\lesssim 0.3\,{\rm dex}$. We comment on this further in §4.2.
@@ 3.2 Weights and Models
where $\rho$ is the probability density function of galaxies that was presented previously in this section, and $W_{i}$ are model-dependent weights that we discuss shortly. Beforehand, we note that our definition above intentionally uses $\rho\left(M_{i},{\rm SFR}_{i}|z_{i}\right)$ rather than $\rho\left(M_{i},{\rm SFR}_{i},z_{i}\right)$. That is, we evaluate the probability of each galaxy being drawn from the underlying distribution given its measured redshift. Stated another way, we avoid weighting galaxies based on their redshift, and only treat the stellar mass and SFR as independent variables. This is because the redshift of our FRB host-galaxy sample is almost certainly biased. In principle this restriction can be lifted if dealing with an unbiased volume-limited galaxy sample and redshift evolution is an important component of models that one wants to test. However we avoid this approach in our present work given limitations of current FRB host-galaxy samples.333 Note that such biases cannot be easily corrected, and extend beyond host-galaxy magnitude selection effects (which we do account for, following K. Sharma et al. 2024). This is because the FRB host-galaxy sample depends also on the radio selection function of FRBs. Given uncertainties about the FRB luminosity function, repetition rate, etc., such effects cannot be easily quantified at present.
@@
[…]
Equation (6) allows for a linear combination of dependence on galaxy stellar mass and SFR. For $B=0$ this reduces to a mass-weighted distribution (Equation 4), as might be expected if—for example—FRBs could be produced by any kind of star and were therefore more likely to occur in galaxies that have a larger number of stars. Conversely, for $A=0$ Equation (6) reduces to a SFR-weighted distribution (Equation 5). This might be expected if FRBs, like CCSNe, were produced only by young stellar populations. Note that the absolute values of the constants $A$ and $B$ are only important if one wants to determine an overall volumetric rate. We are not interested in this question in our present work, and therefore only the ratio of $A/B$ will be important.

### v2

@@ Abstract
We develop a new statistical framework for studying the host galaxies of astrophysical sources that accounts for both redshift evolution and the multi-variate nature of host-galaxy properties. These aspects are critical when dealing with sources that span a wide range of redshifts, and/or with unknown redshift-dependent selection effects. We apply our method to a sample of Fast Radio Burst (FRB) host-galaxies as a means of probing the uncertain progenitor(s) of these events. Using our method we are able to rule out that FRBs track star-formation rate (SFR), as would be expected if FRBs are associated exclusively with young neutron stars born via core-collapse supernovae (SNe). Furthermore, we rule out a recently proposed metallicity-dependent model whereby FRBs track SFR only above an oxygen abundance of $12+\log({\rm O}/{\rm H})\sim 8$. Motivated by the fact that at least one FRB has been localized to a globular cluster (GC), we also investigate the hypothesis that FRB sources track GC mass and explicitly rule out this scenario. Alternatively, we find that a ‘mixed’ model whereby FRBs track a linear combination of both SFR and stellar-mass best explains the data. The preferred parameters of such a mixed model are similar to those inferred for Type Ia SNe, and implies a possible connection between the progenitors of these different transients.
@@
software: numpy (C. R. Harris et al., 2020) matplotlib (J. D. Hunter, 2007), emcee (D. Foreman-Mackey et al., 2013), astropy (Astropy Collaboration et al., 2013, 2018, 2022), scipy (P. Virtanen et al., 2020), pandas (Wes McKinney, 2010), statsmodels (S. Seabold & J. Perktold, 2010)
@@ 1 Introduction
One challenge to the naive magnetar scenario is the environment of some FRB sources. With the growing sample of localized FRBs whose host galaxies have been identified, it has been found that some FRBs reside in quiescent galaxies with very little recent star-formation (K. W. Bannister et al., 2019; V. Ravi et al., 2019; L. Connor et al., 2023; C. J. Law et al., 2024; V. Shah et al., 2025; T. Eftekhari et al., 2025). This is in tension with the conventional expectation for magnetars—that these sources should be formed through the core-collapse of massive stars. In this scenario, FRB sources should closely-track star-formation and their occurrence rate in quiescent galaxies should be fairly low. [^1] Although, as shown by N. Loudas et al. (2025), a small fraction of FRBs would be expected to reside in quiescent galaxies even in scenario where FRBs track star-formation (see also A. E. Nugent et al., 2025, for similar results in the context of core-collapse SNe). More direct evidence along this line has recently surfaced with the localization of a nearby FRB to a globular cluster (GC; F. Kirsten et al. 2022). Redshift zero GCs are characterized by old stellar populations with no ongoing star-formation, directly challenging standard magnetar formation channels. In light of these observations alternative formation channels for magnetar FRB sources have been proposed, including the possibility that magnetars could be formed in the accretion-induced-collapse (AIC) of a white dwarf (WD), or in the aftermath of a subset of binary NS mergers (B. Margalit et al., 2019; K. Kremer et al., 2021; W. Lu et al., 2022; T. Eftekhari et al., 2025). Questions pertaining to the sources of FRBs, and whether some or all of them are related to magnetars (and if so, what are the formation channels and characteristics of such magnetars) are therefore central to our understanding of FRB phenomena.
@@
A particularly powerful tool in addressing these questions is studying the population of host-galaxies and local environments of FRB sources (S. P. Tendulkar et al., 2017; K. W. Bannister et al., 2019; C. D. Bochenek et al., 2020; K. E. Heintz et al., 2020; A. G. Mannings et al., 2021; S. Bhandari et al., 2022; A. C. Gordon et al., 2023; M. Bhardwaj et al., 2024; M. N. Woodland et al., 2024; K. Sharma et al., 2024; N. Loudas et al., 2025; A. C. Gordon et al., 2025; V. Shah et al., 2025; T. Eftekhari et al., 2025). Although host-galaxy studies in general offer only circumstantial evidence about the nature of sources, the ability to statistically study large samples via this approach makes this method appealing. Indeed, host-galaxy studies have been central to building an understanding of other transient phenomena. They have helped solidify the association of (at least typical) long gamma-ray bursts (GRBs) with low-metallicity star-forming environments, consistent with expectations from the collapsar model (A. S. Fruchter et al., 2006); they have shown that superluminous supernovae (SLSNe) occur in similar environments to long GRBs, potentially consistent with rapidly-rotating magnetar sources (R. Lunnan et al., 2014); and they have shown that the formation channels of short GRB and Type Ia supernovae (SNe) progenitors differ significantly from other transients, tracking instead a combination of both star-formation and stellar mass (e.g., E. Scannapieco & L. Bildsten, 2005; W. Fong et al., 2013; A. E. Nugent et al., 2022, 2025).
@@ 2 Past Work
In contrast, host-galaxy properties are intrinsically multivariate and characterized by a combination of several parameters, such as $M$, SFR, and redshift (at the very least). To use KS or AD tests, host-galaxy samples are therefore typically “marginalized” over a number of variables so that the distribution of some single variable (say stellar mass) may be examined. In addition to the convenience of being able to use 1D KS tests, such marginalization also has merit in cases where some variables are more robust than others (indeed, recovering SFRs is significantly more challenging than correctly recovering stellar masses). This approach, though useful, is not statistically sound.
For example, it is possible for two intrinsically different galaxy distributions in the $(M,{\rm SFR})$ plane to have similar 1D distributions in $M$ and/or SFR. Univariate KS or AD tests might therefore suggest that the two samples are drawn from the same distribution, when in fact they are not. One way to conceptualize this ambiguity is to consider that the choice of axes along which to marginalize is arbitrary in a multivariate phase-space. In particular, there is no apriori reason why KS tests should be run on the masses and SFRs of galaxies rather than on some alternative linear combination of the two. The fact that this arbitrary choice of axes will affect the KS score and, perhaps, the resulting conclusion is therefore troubling. Unfortunately, there is no distribution-free generalization of the KS or AD tests to more than one dimension (although see e.g., J. A. Peacock 1983 and G. Fasano & A. Franceschini 1987 for KS generalizations that are only weakly-dependent on the assumed distribution).222‘Distribution-free’ means that there is no statistical method to test whether two multivariate samples are drawn from the same underlying distribution, without making any assumptions about the functional form of the underlying distribution.
@@
This was first pointed out in the context of FRBs by C. D. Bochenek et al. (2021). In that work, the authors devised a method of ‘transposing’ host galaxies to redshift $z=0$. They then compared the transposed host galaxy distribution with a complete sample of core-collapse SN (CCSN) hosts at $z=0$ and found that the two samples are consistent with one another, but only after accounting for this redshift scaling (transposition). This illustrated the importance of accounting for redshift when making such comparisons, which can affect the overall conclusions markedly. However the particular methodology presented by C. D. Bochenek et al. (2021)—though novel and important—has one noticeable shortcoming. In that work, the stellar mass and SFR of each FRB host-galaxy were rescaled to $z=0$ based on the star-forming main sequence of galaxies. Specifically, each galaxy’s stellar-mass and SFR were transposed in a way that preserves the galaxy’s percentiles in both $M$ and SFR relative to the distribution of star-forming galaxies at the redshift of the transient and at $z=0$. Implicit in this choice is the assumption that FRB hosts evolve according to the redshift evolution of star-forming galaxies. However, there is no guarantee apriori that this be the case. Indeed, if one is interested in testing the hypothesis that FRBs track anything other than SFR, this choice may bias the results. An alternative method that avoids ‘transposition’ is presented by K. Sharma et al. (2024) (see their Equation 11). The authors choose a specific property of host galaxies and define a test statistics that evaluates how extreme an FRB host galaxy is compared to the background population at its redshift. While sidestepping the need for ‘transposing’ galaxies to redshift $z=0$, this method operates on only one property at a time, losing potential information encoded in the multivariate nature of host galaxies.
@@ 3.1 Extending the Distribution Function
Being trained on galaxy catalogs, the J. Leja et al. (2022) distribution function is limited by these catalog’s range and completeness. In particular the distribution is valid between redshifts $0.2<z<3$, and is only complete above a redshift-dependent mass-completeness limit $M_{\rm c}(z)$. The latter is due to lower mass galaxies being fainter and harder to detect in surveys. Of course, even though galaxy catalogs may not be complete to such galaxies, galaxies of mass $M<M_{\rm c}(z)$ are in reality abundant. In the FRB host-galaxy sample that we adopt there are two sources (out of 51) that fall below the threshold mass described above. To take these two galaxies into account, and to allow our formalism to be applied more broadly without limitation on galaxy masses, we extend the J. Leja et al. (2022) distribution below $M_{\rm c}$ using the following two considerations.
@@
Here $b(z)$ is the (redshift-dependent) low-mass slope of the star-forming sequence as given by J. Leja et al. (2022), [^3] The parameter $b$ is defined in Equation 9 of that paper. Its dependence on redshift is parameterized via their Equation 10, with best-fit coefficients listed in their Table 1. Here we adopt parameters as fit to the ‘ridge’ of the star-forming sequence. and $\Phi(M,z)$ is a double Schechter-function fit to the stellar mass function taken from J. Leja et al. (2020). Finally, we take $M_{\rm th}=2M_{\rm c}(z)$ where $M_{\rm c}(z)$ is the mass-completeness cutoff described in J. Leja et al. (2022) (e.g., their Table 1). We take a threshold mass that is a factor of two larger than $M_{\rm c}$ to avoid possible edge-effects that may affect the distribution near the mass-completeness cutoff.
@@
The extension described in Equation (3.1) effectively assumes that low mass galaxies with $M<M_{\rm th}$ follow the star-forming sequence (or more accurately, that they can be transposed with relation to the star-forming sequence, even if they fall far from its locus). In general, this is a strong assumption. Many galaxies are quiescent or transitioning and do not follow the star-forming sequence. Given that we are interested in testing the hypothesis that FRB hosts track SFR and/or stellar mass, it is crucial to account for these quiescent galaxies. Indeed, our present work is motivated by shortcomings of previous studies in this respect. Ideally we would not need to extend the distribution function in this way. However we must define some extension if we wish to include the two FRB hosts that fall below the mass-completeness threshold. We argue that Equation (3.1) is nevertheless a reasonable choice for the following reasons. First, low-mass galaxies—those that are affected by our extension—have a strong preference towards being star-forming, while quiescent galaxies (which do not track the star-forming sequence) mostly have $M>M_{\rm th}$. In other words, using the star-forming sequence is a reasonable approximation for low mass galaxies, exactly the regime where Equation (3.1) is relevant. Second, there are only two sources in our FRB sample that are affected by this extension, and both lie only slightly below the cutoff mass. That is, the extrapolation introduced by our extension is very small, and only applies to a small fraction of our sample (see Table 2).
@@
One final extension we apply to the J. Leja et al. (2022) distribution function is a treatment that incorporates low-redshift $z<0.2$ galaxies, where the J. Leja et al. (2022) distribution is truncated. Ideally, a model that is trained on low-redshift galaxy catalogs would be incorporated within this framework and allow a self-consistent evolution with redshift. However this is well beyond the scope of our present work. In order to include low-redshift FRB host galaxies in our current analysis we therefore extend the J. Leja et al. (2022) distribution in redshift-space using the following approximation: we use the extrapolated $z<0.2$ redshift dependence implied by fitting functions to the star-forming sequence and the stellar-mass function obtained by J. Leja et al. (2019, 2022). These are extrapolations because the fits were obtained on data at $z>0.2$. In practice, this implies the following extension,
$\rho\left(M,\,{\rm SFR}\,|\,z<z_{\rm th}\right)=\rho$ $\left[M,\,\frac{\langle{\rm SFR}\rangle(M,z_{\rm th})}{\langle{\rm SFR}\rangle(M,z)}{\rm SFR}\,|\,z\right]$
$\times{\Phi(M,z)}/{\Phi(M,z_{\rm th})}.$ (2)
@@
Here $\langle{\rm SFR}\rangle(M,z)$ is the ‘ridge’ of the star-forming main sequence, as modeled by J. Leja et al. (2022) (see their Equation 9). In Equation (3.1) we take $z_{\rm th}=0.25$ as a cutoff redshift instead of $z=0.2$ in order to avoid possible edge-effects of the distribution near the minimal redshift.
@@
Our redshift extension may be overly simplified, however it is born out of necessity given that we wish to consider FRB host galaxies that would otherwise be excluded from our analysis given the redshift cutoff. Overall, our approach to extending the J. Leja et al. (2022) distribution function to low redshifts and low galaxy masses is very similar to the approach recently taken by K. Sharma et al. (2024) and N. Loudas et al. (2025).
@@ 3.2 Weights and Models
where $\rho$ is the probability density function of galaxies that was presented previously in this section, and $W_{i}$ are model-dependent weights that we discuss shortly. Beforehand, we note that our definition above intentionally uses $\rho\left(M_{i},{\rm SFR}_{i}|z_{i}\right)$ rather than $\rho\left(M_{i},{\rm SFR}_{i},z_{i}\right)$. That is, we evaluate the probability of each galaxy being drawn from the underlying distribution given its measured redshift. Stated another way, we avoid weighting galaxies based on their redshift, and only treat the stellar mass and SFR as independent variables. This is because the redshift of our FRB host-galaxy sample is almost certainly biased. In principle this restriction can be lifted if dealing with an unbiased volume-limited galaxy sample and redshift evolution is an important component of models that one wants to test. However we avoid this approach in our present work given limitations of current FRB host-galaxy samples. [^4] Note that such biases cannot be easily corrected, and extend beyond host-galaxy magnitude selection effects (which we do account for, following K. Sharma et al. 2024). This is because the FRB host-galaxy sample depends also on the radio selection function of FRBs. Given uncertainties about the FRB luminosity function, repetition rate, etc., such effects cannot be easily quantified at present.
@@
$\text{`mixed':}~~~~~W_{i}=A\left(\frac{M_{i}}{M_{\odot}}\right)+B\left(\frac{{\rm SFR}_{i}}{M_{\odot}/{\rm yr}}\right).$ (6)
Equation (6) allows for a linear combination of dependence on galaxy stellar mass and SFR.
For $A=0$ this reduces to a SFR-weighted distribution (Equation 5). This might be expected if FRBs, like CCSNe, were produced only by young stellar populations. Conversely, for $B=0$ Equation (6) reduces to a mass-weighted distribution (Equation 4) that can be interpreted as a preference for older stellar populations. Weighting by mass alone ($B=0$) can also be viewed as a proxy for weighting by the total number of stars (assuming some initial-mass function). Therefore, another possible interpretation of this scenario would be that FRBs can be produced by any kind of star and are therefore more likely to occur in galaxies that have a larger number of stars. Note that the absolute values of the constants $A$ and $B$ are only important if one wants to determine an overall volumetric rate. We are not interested in this question in our present work, and therefore only the ratio of $A/B$ will be important.